Find the length of JK when J is (-5, 4) and K is (3, 2).
Answer by checkley75(3666)

TO FIND THE LENGTHS OF THE TWO SIDES OF THIS RIGHT TRIANGLE WE FIND THE DISTANCE BETWEEN THE X & Y POINTS.THUS:
3+5=8 FOR THE X DISTANCE &
4-2=2 FOR THE Y DISTANCE. NOW WE NEED TO FIND THE HYPOTENUSE . THUS:
2^2+8^2=H^2
4+64=H^2
H^2=68
H=8.246 WHICH IS THE DISTANCE BETWEEN J & K.